OpenStudy (anonymous):

The line through (5,k) and (3,8) has a slope of 4, How might I find what k is?

OpenStudy (marco26):

\[m=\frac{y2-y1}{x2-x1}\] just substitute

OpenStudy (anonymous):

Lets do the formula y=mx+b we know that the slope is 4 so we can replace m with 4. y=4x+b We can plug in the coordinate that we know (3,8) into the equation 8=4(3)+b 8=12+b -12-12 -4=b y=4x-4 k=4(5)-4 k=20-4 k=16 (5,16) Lets graph and check!

OpenStudy (anonymous):

Actually we can use @marco26's response and check 8-16/3-5 -8/-2 4

OpenStudy (anonymous):

cool, thanks!